Consider a schedule with 3 transactions with 2, 3, and 4 operations respectively. Calculate the possible


number of


c) Serial schedules


d) Non serial schedules

1
Expert's answer
2021-11-19T15:32:28-0500


SOLUTION TO THE ABOVE QUESTION



Total Number of Schedules= (2+3+4)!/(2!*3!*4!) = 1260;
 
Total number of serial schedules = Number of different ways of arranging 3 transactions

Total number of serial schedules = 3!
 = 3*2*1
 = 6

Total number of non-serial schedules = Total number of schedules – Total number of serial schedules


= 1260 – 6


= 1254